OpenStudy (anonymous):

If f(x) = x2 - x + 2, then find the value for f(-3).

OpenStudy (anonymous):

plug in "-3" for your "x" values. that's what it pretty much means. Your finding what the function equals to when "x=-3"

OpenStudy (anonymous):

Sorry But I need Show All of my work to get my grade

OpenStudy (anonymous):

Okay. Show your work here then. I'll tell you if you did it right or not. Then just copy what you did here onto your paper.

OpenStudy (anonymous):

No i mean I need The answer with all the steps

OpenStudy (anonymous):

I'm not going to give you the answer without you putting out some work so you can learn also.

OpenStudy (anonymous):

Not whats done here. We help solve the problem so you learn how to do it. He told you how it is done plug in -3 for all your X values and solve. We will let you know if you are right.

OpenStudy (anonymous):

Open Study is the name of the site, not Open Answers

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(-3)=-3 2 -3+2

OpenStudy (anonymous):

Is this correct^^^^^????

OpenStudy (anonymous):

No. Pay attention to your addition and subtractions. You're plugging in a negative 3 not just a regular positive 3. so add them in using parenthesis to make it more clear for you to see. Use the "^" carrot to express exponents.

OpenStudy (anonymous):

for example, say my x = - 4 then my equation would look as so: \[ f(-4) = (-4)^2 - (-4)3 + 2\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[f(-3) = (-3)^2-(-3)+2\]

OpenStudy (anonymous):

f(-3)=(-3)^2-(-3)3+2 Like This???

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

remove the 3 after -(-3)

OpenStudy (anonymous):

now solve it, and you will have your answer. Be careful of the signs. Post answer when done here and someone will make sure you are correct

OpenStudy (anonymous):

I can't solve it I don't know why ??? just can't And if I don''t solve this and 2 more uestions I will be kicked out plzz tell me plzz

OpenStudy (anonymous):

where are you stuck?

OpenStudy (anonymous):

I don't even know where to start?

OpenStudy (anonymous):

You're just adding, subtracting, mutpliying, and taking a constant to the 2nd power. Are you telling us that you don't know how to do that?

OpenStudy (anonymous):

If so, let's start by going term by term.

OpenStudy (anonymous):

I think The answer is 6 but I 'm not sure

OpenStudy (anonymous):

Okay. well lets start off with the first term: \[(-3)^2\] what does this mean? do you know?

OpenStudy (anonymous):

I think we need to multiply -3 times 2

OpenStudy (anonymous):

No. When you see a smaller script that is above a number or letter, that means we times that number or letter by it self by the amount that is displayed, for example, say I have:\[(-2)^2\] it pretty much means\[(-2) \times (-2)\] or if I have, \[(-4)^2\] it means:\[(-4) \times (-4)\] or if the top number is actually 3 instead of 2:\[(-4)^3\] it means:\[(-4) \times (-4) \times (-4)\] do you understand so far?

OpenStudy (anonymous):

okay yes i understand this

OpenStudy (anonymous):

okay, then what does \[(-3)^2\]equal to?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Correct! We solved the first term, now we're subtracting a -3 as so: \[9-(-3)\] what does that equal to?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

now we're adding a positive 2 for the last and final term \[12 +2\]

OpenStudy (anonymous):

what does that equal to?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

is 14 correct??

OpenStudy (anonymous):

YUP YUP! You have just solve each term one by one and have found the answer to your solution. Look at your original equation here:\[f(−3)=(−3)^2−(−3)+2\] you solve the first term so as so: \[f(−3)=9−(−3)+2\]then you solved the next term as so:\[f(−3)=12+2\] then you added the last term\[f(−3)=14\]

OpenStudy (anonymous):

Thank Yuo very much

OpenStudy (anonymous):

You're welcome. Hopefully you have learned that if you get stuck in this type of equation again. Solve them one by one and take your time until you get it so you can do it faster next time.